OpenStudy (anonymous):

what is the cubic function of (1, -20), (7, -20), (2, -40), and (6, -40)

OpenStudy (tkhunny):

ax^3 + bx^2 + cx + d = y (1,-20) a(1)^3 + b(1)^2 + c(1) + d = (-20) (7,-20) a(7)^3 + b(7)^2 + c(7) + d = (-20) (2,-40) a(2)^3 + b(2)^2 + c(2) + d = (-40) (6,-40) a(6)^3 + b(6)^2 + c(6) + d = (-40) You have some solving to do.

OpenStudy (anonymous):

I'm dyslexic I don't understand the numbers mixed with letters

OpenStudy (tkhunny):

Well, then. You're in a heap of trouble in algebra. How did you get this far?

OpenStudy (anonymous):

Don't know

OpenStudy (tkhunny):

You're going to have to provide a hint of some kind. Shall we use the brute-force algebra demonstrated above? What class are you in? What section are you studying?